OpenStudy (anonymous):

use the formula h=vt-16t^2 to find out how long it takes an object to reach a height of 200ft, if it is shot upward with an initial velocity v of 128ft/sec round to the nearest hundredth

OpenStudy (ghazi):

it's H= ut- 1/2 g*t^2 therefore 200= 128t - 5* t^2 , now solve for t

OpenStudy (ghazi):

according to your formula it's 200=128t-16*t^2 which gives 12.5=8t-t^2

OpenStudy (anonymous):

I don't understand how u got that

OpenStudy (ghazi):

i just took out 16 common from the whole equation

OpenStudy (anonymous):

Yes I get that but then what?

OpenStudy (ghazi):

see you can't factorize it so use D=\[(-b \pm \sqrt{b ^{2}-4ac})/2a just find out the discriminant of quadratic equation \]

OpenStudy (ghazi):

find out the discriminant of quadratic equation

OpenStudy (anonymous):

so I would bring the 12.5 over and use the quadratic equation?

OpenStudy (ghazi):

yes just take all the elements of right side to the left and your equation would be t^2-8t+12.5=0 i hope now you can do it